# Learning multi-modal generative models with permutation-invariant encoders and tighter variational bounds

### v2

@@ Appendix D Permutation-invariant architectures
We introduce here a standard multi-head attention (Bahdanau et al., 2014; Vaswani et al., 2017) mapping $\text{MHA}_{\vartheta}\colon\mathbb{R}^{I\times D_{X}}\times\mathbb{R}^{S\times D_{Y}}\to\mathbb{R}^{I\times D_{Y}}$ given by
@@
where $\mu$ and $\sigma$ standardise the input $h$ by computing the mean, and the variance, respectively, over some axis of $h$, whilst $\gamma$ and $\beta$ define a transformation. LayerNorm (Ba et al., 2016) standardises inputs over the last axis, e.g., $\mu(h)=\frac{1}{D}\sum_{d=1}^{D}\mu_{\cdot,d}$, i.e., separately for each element. In contrast, SetNorm (Zhang et al., 2022b) standardises inputs over both axes, e.g., $\mu(h)=\frac{1}{TD}\sum_{t=1}^{T}\sum_{d=1}^{D}\mu_{t,d}$, thereby losing the global mean and variance only. In both cases, $\gamma$ and $\beta$ share their values across the first axis. Both normalizations are permutation-equivariant.
@@
We consider a masked pre-layer-norm (Wang et al., 2019a; Xiong et al., 2020) multi-head transformer block
@@
We have introduced a multi-modal generative model in a general form that also applies to the time-series setup, such as when a latent Markov process drives multiple time series. For example, consider a latent Markov process $Z=(Z_{t})_{t\in\mathbb{N}}$ with prior dynamics $p_{\theta}(z_{1},\ldots,z_{T})=p_{\theta}(z_{1})\prod_{t=2}^{T}p_{\theta}(z_{t}|z_{t-1})$ for an initial density $p_{\theta}(z_{1})$ and homogeneous Markov kernels $p_{\theta}(z_{t}|z_{t-1})$. Conditional on $Z$, suppose that the time-series $(X_{s,t})_{t\in\mathbb{N}}$ follows the dynamics $p_{\theta}(x_{s,1},\ldots,x_{s,T}|z_{1},\ldots,z_{T})=\prod_{t=2}^{T}p_{\theta}(x_{s,t}|z_{t})$ for decoding densities $p_{\theta}(x_{s,t}|z_{t})$. A common choice (Chung et al., 2015) for modeling the encoding distribution for such sequential (uni-modal) VAEs is to assume the factorisation $q_{\phi}(z_{1},\ldots z_{T}|x_{1},\ldots x_{T})=q_{\phi}(z_{1}|x_{1})\prod_{t=2}^{T}q_{\phi}(z_{t}|z_{t-1},x_{t})$ for $x_{t}=(x_{s,t})_{s\in\mathcal{M}}$, with initial encoding densities $q_{\phi}(z_{1}|x_{1})$ and encoding Markov kernels $q_{\phi}(z_{t}|z_{t-1},x_{t})$. One can again consider modality-specific encodings $h_{s}=(h_{s,1},\ldots,h_{s,T})$, $h_{s,t}=h_{s,\varphi}(x_{s,t})$, now applied separately at each time step that are then used to construct Markov kernels that are permutation-invariant in the form of $q^{\prime}_{\phi}(z_{t}|z_{t-1},\pi h_{\varphi}(x_{t,\mathcal{S}}))=q^{\prime}_{\phi}(z_{t}|z_{t-1},h_{\varphi}(x_{t,\mathcal{S}}))$ for permutations $\pi\in\mathbb{S}_{\mathcal{S}}$. Alternatively, in the absence of the auto-regressive encoding structure with Markov kernels, one could also use transformer models that use absolute or relative positional embeddings across the last temporal axis but no positional embeddings across the first modality axis, followed by a sum-pooling operation across the modality axis. Note that previous works using multi-modal time series such as Kramer et al. (2022) use a non-amortized encoding distribution for the full multi-modal posterior only. A numerical evaluation of permutation-invariant schemes for time series models is, however, outside the scope of this work.
@@ Appendix E Permutation-equivariance and private latent variables
To compute the multi-modal variational bounds, notice that the required KL-divergences can be written as follows:
$\mathsf{K}\mathsf{L}(q_{\phi}(z^{\prime},\tilde{z}|x_{\mathcal{S}})|p_{\theta}(z^{\prime},\tilde{z}))=\mathsf{K}\mathsf{L}(q_{\phi}(z^{\prime}|x_{\mathcal{S}})|p_{\theta}(z^{\prime}))+\int q_{\phi}(z^{\prime}|x_{\mathcal{S}})\mathsf{K}\mathsf{L}(q_{\phi}(\tilde{z}\mathcal{S}|z^{\prime},x_{\mathcal{S}})|p_{\theta}(\tilde{z}_{\mathcal{S}}|z^{\prime}))\mathrm{d}z^{\prime}$
@@
It is instructive to compare our bound with the MMVAE+ approach suggested in Palumbo et al. (2023). Assuming a uniform masking distribution restricted to uni-modal sets so that $\mathcal{S}=\{s\}$ for some $s\in\mathcal{M}$, we can write the bound from Palumbo et al. (2023) as $\frac{1}{M}\sum_{s=1}^{M}\mathcal{L}^{\text{MMVAE+}}_{\{s\}}(x)$ with
@@ Appendix G Mixture model extensions for different variational bounds
We will pursue here an encoding approach that does not require modeling the encoding distribution over the discrete latent variables explicitly, thus avoiding large variances in score-based Monte Carlo estimators (Ranganath et al., 2014) or resorting to advanced variance reduction techniques (Kool et al., 2019) or alternatives such as continuous relaxation approaches (Jang et al., 2016; Maddison et al., 2016).
@@ Appendix H Algorithm and STL-gradient estimators
We consider a multi-modal extension of the sticking-the-landing (STL) gradient estimator (Roeder et al., 2017) that has also been used in previous multi-modal bounds (Shi et al., 2019). The gradient estimator ignores the score function terms when sampling $q_{\phi}(z|x_{\mathcal{S}})$ for variance reduction purposes because it has a zero expectation. For the bounds (2) that involves sampling from $q_{\phi}(z|x_{\mathcal{S}})$ and $q_{\phi}(z|x_{\mathcal{M}})$, we thus ignore the score terms for both integrals. Consider the reparameterization with noise variables $\epsilon_{\mathcal{S}}$, $\epsilon_{\mathcal{M}}\sim p$ and transformations $z_{\mathcal{S}}=t_{{\mathcal{S}}}(\phi,\epsilon_{\mathcal{S}},x_{\mathcal{S}})=f_{\text{invariant-agg}}(\vartheta,\epsilon_{\mathcal{S}},\mathcal{S},h_{\mathcal{S}})$, for $h_{\mathcal{S}}=h_{\varphi,s}(x_{s})_{s\in\mathcal{S}}$ and $z_{\mathcal{M}}=t_{\mathcal{M}}(\phi,\epsilon_{\mathcal{M}},x_{\mathcal{M}})=f_{\text{invariant-agg}}(\vartheta,\epsilon_{\mathcal{M}},\mathcal{M},h_{\mathcal{M}})$, for $h_{\mathcal{M}}=h_{\varphi,s}(x_{s})_{s\in\mathcal{M}}$ We need to learn only a single aggregation function that applies that masks the modalities appropriately. Pseudo-code for computing the gradients are given in Algorithm 1. If the encoding distribution is a mixture distribution, we apply the stop-gradient operation also to the mixture weights. Notice that in the case of a mixture prior and an encoding distribution that includes the mixture component, the optimal encoding density over the mixture variable has no variational parameters and is given as the posterior density of the mixture component under the generative parameters of the prior.
@@ Appendix I Evaluation of multi-modal generative models
We evaluate models using different metrics suggested previously for multi-modal learning, see for example Shi et al. (2019); Wu and Goodman (2019); Sutter et al. (2021).
@@
Following previous work (Shi et al., 2019; Sutter et al., 2021; Daunhawer et al., 2022; Javaloy et al., 2022), we assess whether the generated data share the same information in the form of the class labels across different modalities. To do so, we use pre-trained classifiers $\text{clf}_{s}\colon\mathsf{X}_{s}\to[K]$ that classify values from modality $s$ to $K$ possible classes. More precisely, for $\mathcal{S}\subset\mathcal{M}$ and $m\in\mathcal{M}$, we compute the self- ($m\in\mathcal{S}$) or cross- ($m\notin\mathcal{S}$) coherence $\text{C}_{\mathcal{S}\to m}$ as the empirical average of
@@
Here, $W^{\prime}_{s},\tilde{W}_{s}\in\mathbb{R}^{D_{s}\times D^{\prime}}$ with $D^{\prime}=D/(M+1)=5$, Furthermore, the latent variable $Z$ can be written as $Z=(Z^{\prime},\tilde{Z}_{1},\ldots,\tilde{Z}_{M})$ for private and shared latent variables $\tilde{Z}_{s}$, resp. $Z^{\prime}$. We similarly generate orthonormal $\begin{bmatrix}W^{\prime}_{s},\tilde{W}_{s}\end{bmatrix}$ from a QR decomposition. Observe that the general generative model with latent variable $Z$ corresponds to the generative model (8) with shared $Z^{\prime}$ and private latent variables $\tilde{Z}$ with straightforward adjustments for the decoding functions. Similar models have been considered previously, particularly from a Bayesian standpoint with different sparsity assumptions on the generative parameters (Archambeau and Bach, 2008; Virtanen et al., 2012; Zhao et al., 2016).
@@
Figure 4: Bi-modal non-linear model with label and continuous modality based on our bound. SumP: SumPooling, SumPM: SumPoolingMixture.
@@
The MNIST-SVHN-Text data set is taken from the code accompanying Sutter et al. (2021) with around 1.1 million train and 200k test samples. All models are trained for 100 epochs with a batch size of 250 using Adam (Kingma and Ba, 2014) and a cosine decay schedule from 0.0005 to 0.0001.
@@
Our bound Mixture bound
@@
Results from Sutter et al. (2021) and Sutter et al. (2020)
@@
Our bound Mixture bound
@@
Results from Sutter et al. (2021), Sutter et al. (2020) and Hwang et al. (2021)
@@
Our bound Mixture bound
@@
Our bound Mixture bound
@@
Our bound Mixture bound
@@
Our bound Mixture bound
@@
Latent classification accuracy.
@@
Our bound Mixture bound
@@
Results from Sutter et al. (2021), Sutter et al. (2020) and Hwang et al. (2021)
@@
Our bound Mixture bound
@@
For SVHN and Text, we use 2d- or 1d-convolutional layers, respectively, denoted as Conv($f,k,s$) for feature dimension $f$, kernel-size $k$, and stride $s$. We denote transposed convolutions as tConv. We use the neural network architectures as implemented in Flax Heek et al. (2023).
@@
(b) SVHN-specific encoding functions $h_{s}(x_{s})$. Modality dimensions $D_{s}=3\times 32\times 32$. Embedding dimension is $D_{E}=2D$ for PoE/MoE and $D_{E}=256$ for SumPooling/SelfAttention. For PoE+/MoE+, we add four times a Dense layer of size $256$ with ReLU layer before the last linear layer.
@@
M.6 MNIST-SVHN-Text with private latent variables
@@
For models with private latent variables, we concatenate the shared and private latent variables. We use a Laplace likelihood as the decoding distribution for MNIST and SVHN, where the decoder function learns both its mean as a function of the latent and a constant log-standard-deviation at each pixel. Following previous works (Shi et al., 2019; Sutter et al., 2021), we re-weight the log-likelihoods for different modalities relative to their dimensions.
@@
Our computations were performed on shared HPC systems. All experiments except Section 5.3 were run on a CPU server using one or two CPU cores. The experiments in Section 5.3 were run on a GPU server using one NVIDIA A100.
Our implementation is based on JAX (Bradbury et al., 2018) and Flax (Heek et al., 2023). We compute the mean correlation co-efficient (MCC) between true and inferred latent variables following Khemakhem et al. (2020b), as in https://github.com/ilkhem/icebeem and follow the data and model generation from Khemakhem et al. (2020a), https://github.com/ilkhem/iVAE in Section 5.2, as well as https://github.com/hanmenghan/CPM_Nets from Zhang et al. (2019) for generating the missingness mechanism. In our MNIST-SVHN-Text experiments, we use code from Sutter et al. (2021), https://github.com/thomassutter/MoPoE.

### v3

@@ Appendix D Permutation-invariant architectures
We introduce here a standard multi-head attention [Bahdanau et al., 2014, Vaswani et al., 2017] mapping $\text{MHA}_{\vartheta}\colon\mathbb{R}^{I\times D_{X}}\times\mathbb{R}^{S\times D_{Y}}\to\mathbb{R}^{I\times D_{Y}}$ given by
@@
where $\mu$ and $\sigma$ standardize the input $h$ by computing the mean, and the variance, respectively, over some axis of $h$, whilst $\gamma$ and $\beta$ define a transformation. LayerNorm [Ba et al., 2016] standardises inputs over the last axis, e.g., $\mu(h)=\frac{1}{D}\sum_{d=1}^{D}\mu_{\cdot,d}$, i.e., separately for each element. In contrast, SetNorm [Zhang et al., 2022b] standardises inputs over both axes, e.g., $\mu(h)=\frac{1}{TD}\sum_{t=1}^{T}\sum_{d=1}^{D}\mu_{t,d}$, thereby losing the global mean and variance only. In both cases, $\gamma$ and $\beta$ share their values across the first axis. Both normalizations are permutation-equivariant.
@@
We consider a masked pre-layer-norm [Wang et al., 2019a, Xiong et al., 2020] multi-head transformer block
@@
Remark 20 (Context-aware pooling)
$o_{s}=\frac{1}{Z}\sum_{t\in\mathcal{S}}\kappa(g_{s},g_{t})v(g_{t}),$
can be seen as a learnable non-symmetric kernel [Wright and Gonzalez, 2021, Cao, 2021]. Conceptually, the attention encoder pools a learnable $D$-dimensional function $v$ using a learnable context-dependent weighting function. While such attention models directly account for the interaction between the different encodings, a DeepSet aggregation approach may require a sufficiently high-dimensional latent space $D_{P}$ to achieve universal approximation properties [Wagstaff et al., 2022].
@@
We have introduced a multi-modal generative model in a general form that also applies to the time-series setup, such as when a latent Markov process drives multiple time series. For example, consider a latent Markov process $Z=(Z_{t})_{t\in\mathbb{N}}$ with prior dynamics $p_{\theta}(z_{1},\ldots,z_{T})=p_{\theta}(z_{1})\prod_{t=2}^{T}p_{\theta}(z_{t}|z_{t-1})$ for an initial density $p_{\theta}(z_{1})$ and homogeneous Markov kernels $p_{\theta}(z_{t}|z_{t-1})$. Conditional on $Z$, suppose that the time-series $(X_{s,t})_{t\in\mathbb{N}}$ follows the dynamics $p_{\theta}(x_{s,1},\ldots,x_{s,T}|z_{1},\ldots,z_{T})=\prod_{t=2}^{T}p_{\theta}(x_{s,t}|z_{t})$ for decoding densities $p_{\theta}(x_{s,t}|z_{t})$. A common choice [Chung et al., 2015] for modeling the encoding distribution for such sequential (uni-modal) VAEs is to assume the factorization $q_{\phi}(z_{1},\ldots z_{T}|x_{1},\ldots x_{T})=q_{\phi}(z_{1}|x_{1})\prod_{t=2}^{T}q_{\phi}(z_{t}|z_{t-1},x_{t})$ for $x_{t}=(x_{s,t})_{s\in\mathcal{M}}$, with initial encoding densities $q_{\phi}(z_{1}|x_{1})$ and encoding Markov kernels $q_{\phi}(z_{t}|z_{t-1},x_{t})$. One can again consider modality-specific encodings $h_{s}=(h_{s,1},\ldots,h_{s,T})$, $h_{s,t}=h_{s,\varphi}(x_{s,t})$, now applied separately at each time step that are then used to construct Markov kernels that are permutation-invariant in the form of $q^{\prime}_{\phi}(z_{t}|z_{t-1},\pi h_{\varphi}(x_{t,\mathcal{S}}))=q^{\prime}_{\phi}(z_{t}|z_{t-1},h_{\varphi}(x_{t,\mathcal{S}}))$ for permutations $\pi\in\mathbb{S}_{\mathcal{S}}$. Alternatively, in the absence of the auto-regressive encoding structure with Markov kernels, one could also use transformer models that use absolute or relative positional embeddings across the last temporal axis but no positional embeddings across the first modality axis, followed by a sum-pooling operation across the modality axis. Note that previous works using multi-modal time series such as Kramer et al. [2022] use a non-amortized encoding distribution for the full multi-modal posterior only. A numerical evaluation of permutation-invariant schemes for time series models is, however, outside the scope of this work.
Remark 22 (Alternative multi-modal encoding models)
Learning different encoders for each modality subset can be an alternative when the number of modalities is small. For example, one could learn different MLP heads for each modality subset $\mathcal{S}$ that aggregates the encoded features $h_{\mathcal{S}}$ from modality-specific encoders that are shared for any modality mask. Our initial experiments with such encoders for the simulated linear modalities in Section 5.1 for $M=5$ modalities did not improve on the permutation-invariant models, while also being more computationally demanding.
@@ Appendix E Permutation-equivariance and private latent variables
To compute the multi-modal variational bounds, notice that the required KL divergences can be written as follows:
$\mathsf{K}\mathsf{L}(q_{\phi}(z^{\prime},\tilde{z}|x_{\mathcal{S}})|p_{\theta}(z^{\prime},\tilde{z}))=\mathsf{K}\mathsf{L}(q_{\phi}(z^{\prime}|x_{\mathcal{S}})|p_{\theta}(z^{\prime}))+\int q_{\phi}(z^{\prime}|x_{\mathcal{S}})\mathsf{K}\mathsf{L}(q_{\phi}(\tilde{z}_{\mathcal{S}}|z^{\prime},x_{\mathcal{S}})|p_{\theta}(\tilde{z}_{\mathcal{S}}|z^{\prime}))\mathrm{d}z^{\prime}$
@@
It is instructive to compare our bound with the MMVAE+ approach suggested in Palumbo et al. [2023]. Assuming a uniform masking distribution restricted to uni-modal sets so that $\mathcal{S}=\{s\}$ for some $s\in\mathcal{M}$, we can write the bound from Palumbo et al. [2023] as $\frac{1}{M}\sum_{s=1}^{M}\mathcal{L}^{\text{MMVAE+}}_{\{s\}}(x)$ with
@@
Remark 25 (Cross-modal context variables and permutation-equivariant models)
In contrast to the PoE model, where the private encodings are independent, the private encodings are dependent in the Sum-Pooling model by conditioning on a sample from the shared latent space. The shared latent variable $Z^{\prime}$ can be seen as a shared cross-modal context variable, and similar probabilistic constructions to encode such context variables via permutation-invariant models have been suggested in few-shot learning algorithms [Edwards and Storkey, 2016, Giannone and Winther, 2022] or, particularly, for neural process models [Garnelo et al., 2018b, a, Kim et al., 2018]. Permutation-equivariant models have been studied for stochastic processes where invariant priors correspond to equivariant posteriors [Holderrieth et al., 2021], such as Gaussian processes or Neural processes with private latent variables, wherein dependencies in the private latent variables can be constructed hierarchically [Wang and Van Hoof, 2020, Xu et al., 2023].
@@ Appendix G Mixture model extensions for different variational bounds
We will pursue here an encoding approach that does not require modeling the encoding distribution over the discrete latent variables explicitly, thus avoiding large variances in score-based Monte Carlo estimators [Ranganath et al., 2014] or resorting to advanced variance reduction techniques [Kool et al., 2019] or alternatives such as continuous relaxation approaches [Jang et al., 2016, Maddison et al., 2016].
@@ Appendix H Algorithm and STL-gradient estimators
We consider a multi-modal extension of the sticking-the-landing (STL) gradient estimator [Roeder et al., 2017] that has also been used in previous multi-modal bounds [Shi et al., 2019]. The gradient estimator ignores the score function terms when sampling $q_{\phi}(z|x_{\mathcal{S}})$ for variance reduction purposes because it has a zero expectation. For the bounds (2) that involves sampling from $q_{\phi}(z|x_{\mathcal{S}})$ and $q_{\phi}(z|x_{\mathcal{M}})$, we thus ignore the score terms for both integrals. Consider the reparameterization with noise variables $\epsilon_{\mathcal{S}}$, $\epsilon_{\mathcal{M}}\sim p$ and transformations $z_{\mathcal{S}}=t_{{\mathcal{S}}}(\phi,\epsilon_{\mathcal{S}},x_{\mathcal{S}})=f_{\text{invariant-agg}}(\vartheta,\epsilon_{\mathcal{S}},\mathcal{S},h_{\mathcal{S}})$, for $h_{\mathcal{S}}=h_{\varphi,s}(x_{s})_{s\in\mathcal{S}}$ and $z_{\mathcal{M}}=t_{\mathcal{M}}(\phi,\epsilon_{\mathcal{M}},x_{\mathcal{M}})=f_{\text{invariant-agg}}(\vartheta,\epsilon_{\mathcal{M}},\mathcal{M},h_{\mathcal{M}})$, for $h_{\mathcal{M}}=h_{\varphi,s}(x_{s})_{s\in\mathcal{M}}$ We need to learn only a single aggregation function that applies and masks the modalities appropriately. Pseudo-code for computing the gradients are given in Algorithm 1. If the encoding distribution is a mixture distribution, we apply the stop-gradient operation also to the mixture weights. Notice that in the case of a mixture prior and an encoding distribution that includes the mixture component, the optimal encoding density over the mixture variable has no variational parameters and is given as the posterior density of the mixture component under the generative parameters of the prior.
@@ Appendix I Evaluation of multi-modal generative models
We evaluate models using different metrics suggested previously for multi-modal learning, see for example Shi et al. [2019], Wu and Goodman [2019], Sutter et al. [2021].
@@
Following previous work [Shi et al., 2019, Sutter et al., 2021, Daunhawer et al., 2022, Javaloy et al., 2022], we assess whether the generated data share the same information in the form of the class labels across different modalities. To do so, we use pre-trained classifiers $\text{clf}_{s}\colon\mathsf{X}_{s}\to[K]$ that classify values from modality $s$ to $K$ possible classes. More precisely, for $\mathcal{S}\subset\mathcal{M}$ and $m\in\mathcal{M}$, we compute the self- ($m\in\mathcal{S}$) or cross- ($m\notin\mathcal{S}$) coherence $\text{C}_{\mathcal{S}\to m}$ as the empirical average of
@@
Here, $W^{\prime}_{s},\tilde{W}_{s}\in\mathbb{R}^{D_{s}\times D^{\prime}}$ with $D^{\prime}=D/(M+1)=5$, Furthermore, the latent variable $Z$ can be written as $Z=(Z^{\prime},\tilde{Z}_{1},\ldots,\tilde{Z}_{M})$ for private and shared latent variables $\tilde{Z}_{s}$, resp. $Z^{\prime}$. We similarly generate orthonormal $\begin{bmatrix}W^{\prime}_{s},\tilde{W}_{s}\end{bmatrix}$ from a QR decomposition. Observe that the general generative model with latent variable $Z$ corresponds to the generative model (9) with shared $Z^{\prime}$ and private latent variables $\tilde{Z}$ with straightforward adjustments for the decoding functions. Similar models have been considered previously, particularly from a Bayesian standpoint with different sparsity assumptions on the generative parameters [Archambeau and Bach, 2008, Virtanen et al., 2012, Zhao et al., 2016].
@@
Figure 6: Bi-modal non-linear model with label and continuous modality based on our proposed objective. SumP: SumPooling, SumPM: SumPoolingMixture.
@@
The MNIST-SVHN-Text data set is taken from the code accompanying Sutter et al. [2021] with around 1.1 million train and 200k test samples. All models are trained for 100 epochs with a batch size of 250 using Adam [Kingma and Ba, 2014] and a cosine decay schedule from 0.0005 to 0.0001.
@@
Proposed objective Mixture bound
@@
Results from Sutter et al. [2021] and Sutter et al. [2020]
@@
Proposed objective Mixture bound
@@
Results from Sutter et al. [2021], Sutter et al. [2020] and Hwang et al. [2021]
@@
Proposed objective Mixture bound
@@
Proposed objective Mixture bound
@@
Proposed objective Mixture bound
@@
Proposed objective Mixture bound
@@
L.6 Latent classification accuracy
@@
Proposed objective Mixture bound
@@
Results from Sutter et al. [2021], Sutter et al. [2020] and Hwang et al. [2021]
@@
Proposed objective Mixture bound
@@
For SVHN and Text, we use 2d- or 1d-convolutional layers, respectively, denoted as Conv($f,k,s$) for feature dimension $f$, kernel-size $k$, and stride $s$. We denote transposed convolutions as tConv. We use the neural network architectures as implemented in Flax Heek et al. [2023].
@@
(b) SVHN-specific encoding functions $h_{s}(x_{s})$. Modality dimensions $D_{s}=3\times 32\times 32$. The embedding dimension is $D_{E}=2D$ for PoE/MoE and $D_{E}=256$ for SumPooling/SelfAttention. For PoE+/MoE+, we add four times a Dense layer of size $256$ with ReLU layer before the last linear layer.
@@
For models with private latent variables, we concatenate the shared and private latent variables. We use a Laplace likelihood as the decoding distribution for MNIST and SVHN, where the decoder function learns both its mean as a function of the latent and a constant log-standard deviation at each pixel. Following previous works [Shi et al., 2019, Sutter et al., 2021], we re-weight the log-likelihoods for different modalities relative to their dimensions.
@@
A reference implementation is available at https://github.com/marcelah/MaskedMultimodalVAE. Our computations were performed on shared HPC systems. All experiments except Section 5.3 were run on a CPU server using one or two CPU cores. The experiments in Section 5.3 were run on a GPU server using one NVIDIA A100.
Our implementation is based on JAX [Bradbury et al., 2018] and Flax [Heek et al., 2023]. We compute the mean correlation coefficient (MCC) between true and inferred latent variables following Khemakhem et al. [2020b], as in https://github.com/ilkhem/icebeem and follow the data and model generation from Khemakhem et al. [2020a], https://github.com/ilkhem/iVAE in Section 5.2, as well as https://github.com/hanmenghan/CPM_Nets from Zhang et al. [2019] for generating the missingness mechanism. In our MNIST-SVHN-Text experiments, we use code from Sutter et al. [2021], https://github.com/thomassutter/MoPoE.